# Distance Between Stochastic Linear Systems

### v1

@@ 5 Comparing Time Domain & Frequency Domain Distance Measures
$\underbrace{\inf_{\pi_{\mathbf{P}}\in\Pi(\mathbf{f}_{P_{1}},\mathbf{f}_{P_{2}})}\mathrm{d}_{q}(P_{1},P_{2})}_{=\mathrm{d}_{q}(P_{1},P_{2})}$ $\leq\inf_{\pi_{\mathbf{P}}\in\Pi(\mathbf{f}_{P_{1}},\mathbf{f}_{P_{2}})}\int_{\mathrm{supp}\left(\mathbf{f}_{P_{1}}\right)\times\mathrm{supp}\left(\mathbf{f}_{P_{2}}\right)}\!\!\delta_{g}(P_{1},P_{2})^{q}\,\pi_{\mathbf{P}}(dP_{1},dP_{2})$
$=\mathrm{dist}_{\Sigma_{1},\Sigma_{2},\delta_{g}}.$
@@
(72) $\bar{P}_{1}(s)=\frac{1}{s^{2}+1.26s+3.24},\quad\bar{P}_{2}(s)=\frac{1}{s^{2}+1.32s+1.44}.$
@@
To compute the frequency domain distance, a frequency grid in the log space between $10^{-2},10^{2}]$ rad/sec was discretized into $M=100$ points. Precisely speaking, $\Omega_{M}=\mathrm{logspace}(10^{-2},10^{2},100)$. The quantities of interests namely the frequency domain distance $\mathrm{d}_{q}(P_{1},P_{2})$ and the time domain distance $\mathrm{dist}_{\Sigma_{1},\Sigma_{2},\delta_{g}}$ were computed using (3.6) and (48) respectively. In both the distance computations, the corresponding type-$1$ Wasserstein distance computation was carried out using the linear programming technique. As a result of the computation, we obtained the frequency domain distance $\mathrm{d}_{q}(P_{1},P_{2})=0.3795$, the time domain distance $\mathrm{dist}_{\Sigma_{1},\Sigma_{2},\delta_{g}}=0.3812$ and the gap metric between the nominal models given by (72) was found out to be $0.3822$. Clearly as expected, we obtained $\mathrm{d}_{q}(P_{1},P_{2})\leq\mathrm{dist}_{\Sigma_{1},\Sigma_{2},\delta_{g}}$ and thereby agreeing to the claims of Theorem˜5.5.

### v2

@@ 5 Comparing Time Domain & Frequency Domain Distance Measures
$\underbrace{\inf_{\pi_{\mathbf{P}}\in\Pi(\mathbf{f}_{P_{1}},\mathbf{f}_{P_{2}})}\mathrm{d}_{q}(P_{1},P_{2})}_{=\mathrm{d}_{q}(P_{1},P_{2})}$
$\leq\inf_{\pi_{\mathbf{P}}\in\Pi(\mathbf{f}_{P_{1}},\mathbf{f}_{P_{2}})}\int_{\mathrm{supp}\left(\mathbf{f}_{P_{1}}\right)\times\mathrm{supp}\left(\mathbf{f}_{P_{2}}\right)}\!\!\delta_{g}(P_{1},P_{2})^{q}\,\pi_{\mathbf{P}}(dP_{1},dP_{2})$
$=\mathrm{dist}^{q}_{\Sigma_{1},\Sigma_{2},\delta_{g}}.$
This completes the proof. ∎
@@
$\bar{P}_{1}(s)=\frac{1}{s^{2}+1.26s+3.24},\quad\bar{P}_{2}(s)=\frac{1}{s^{2}+1.32s+1.44}.$ (73)
@@
To compute the frequency domain distance, a frequency grid in the log space between $10^{-2},10^{2}]$ rad/sec was discretized into $M=100$ points. Precisely speaking, $\Omega_{M}=\mathrm{logspace}(10^{-2},10^{2},100)$. The quantities of interests namely the frequency domain distance $\mathrm{d}_{q}(P_{1},P_{2})$ and the time domain distance $\mathrm{dist}_{\Sigma_{1},\Sigma_{2},\delta_{g}}$ were computed using (3.3) and (49) respectively. In both the distance computations, the corresponding type-$1$ Wasserstein distance computation was carried out using the linear programming technique. As a result of the computation, we obtained the frequency domain distance $\mathrm{d}_{q}(P_{1},P_{2})=0.3795$, the time domain distance $\mathrm{dist}_{\Sigma_{1},\Sigma_{2},\delta_{g}}=0.3812$ and the gap metric between the nominal models given by (73) was found out to be $0.3822$. Clearly as expected, we obtained $\mathrm{d}_{q}(P_{1},P_{2})\leq\mathrm{dist}_{\Sigma_{1},\Sigma_{2},\delta_{g}}$ and thereby agreeing to the claims of Theorem 5.3.